### v1

@@ 5.6.1 Global sensitivity analysis
We aim to quantify how input parameters drive model outputs by applying multiple global sensitivity methods and checking the consistency of their conclusions. Global sensitivity analysis identifies influential inputs, dominant interactions, and how input uncertainty propagates to output variability. Although decision trees are limited and discontinuous, they are sequences of simple procedures that are easy to interpret. For example, the first node of the GB or DT is “intolerance threshold below 0.63”, which confirms what was graphically observed on the complete dataset. In a random forest, the importance of a variable is often measured via the Mean Decrease in Impurity (MDI). At each split of a tree, the reduction in impurity, often measured by the Gini index, quantifies the probability that a sample is misclassified. This reduction is attributed to the variable that enabled the split, and by summing these reductions over all trees, a global measure of the importance of each feature is obtained. However, this approach can bias the estimation in favor of variables with high cardinality, which offer more splitting possibilities. In this way, in our test case, we obtain the importances shown in Fig. 13. For both outputs, the variables that most influence the model’s predictions, as identified on the extended dataset, are clearly found to be intolerance for convergence and density for sparsity. It is also observed that grid size has a rather weak influence on the outputs. Equivalent importance analyses can also be performed with other models. For instance, applying the same procedure to a k-nearest neighbors model and aggregating SHAP absolute values produces the results shown in Fig. 14. The predicted importances differ significantly from those of the random forest analysis as most variables appear significant. This occurs because kNN SHAP values primarily reflect local, instance-specific effects rather than global trends, making them less suitable for traditional global sensitivity analysis. Appendix A further illustrates that importance outcomes vary depending on whether the surrogate captures broad system-level tendencies or local deviations and as such the SHAP importance of a Gaussian process are aligned with the random forest results of Fig. 14.
@@ 6 Conclusions and perspectives
The codes and data used to produce the results are open-source and available online: https://github.com/ANR-MIMICO/EAAI_XAI.
@@ Appendix A Complementary Local Analysis of Variable Importance
In conclusion, SHAP values help explain model predictions, but their interpretation differs depending on the algorithm. Orange models, such as GP, reflect how each feature affects the predicted mean and uncertainty of the output, and SHAP highlights feature importance in shaping the underlying function rather than direct instance similarity. On the contrary, green models such as kNN make predictions based on the similarity between data points, using distances in feature space. SHAP values here indicate how much each feature contributes to determining the nearest neighbors and shaping the final prediction. There is no clear response on what is a “really important” variable, except that it depends on whether the interest of the practitioner is more in the global trends or in the potential local effects. Moreover, studying a complex system through a surrogate model is a strongly impacting method to gather information as most of the resulting information comes from the structure of the surrogate model itself instead of coming from the complex system.

### v2

@@ 5.6.1 Global sensitivity analysis
We aim to quantify how input parameters drive model outputs by applying multiple global sensitivity methods and checking the consistency of their conclusions. Global sensitivity analysis identifies influential inputs, dominant interactions, and how input uncertainty propagates to output variability. Although decision trees are limited and discontinuous, they are sequences of simple procedures that are easy to interpret. For example, the first node of the GB or DT is “intolerance threshold below 0.63”, which confirms what was graphically observed on the complete dataset. In a random forest, the importance of a variable is often measured via the Mean Decrease in Impurity (MDI). At each split of a tree, the reduction in impurity, often measured by the Gini index, quantifies the probability that a sample is misclassified. This reduction is attributed to the variable that enabled the split, and by summing these reductions over all trees, a global measure of the importance of each feature is obtained. However, this approach can bias the estimation in favor of variables with high cardinality, which offer more splitting possibilities. In this way, in our test case, we obtain the importances shown in Fig. 13. For both outputs, the variables that most influence the model’s predictions, as identified on the extended dataset, are clearly found to be intolerance for convergence and density for sparsity. It is also observed that grid size has a rather weak influence on the outputs. Equivalent importance analyses can also be performed with other models. For instance, applying the same procedure to a k-nearest neighbors model and aggregating SHAP absolute values produces the results shown in Fig. 14. The predicted importances differ significantly from those of the random forest analysis, as most variables appear significant. This occurs because kNN SHAP values primarily reflect local, instance-specific effects rather than global trends, making them less suitable for traditional global sensitivity analysis. Appendix A further illustrates that importance outcomes vary depending on whether the surrogate captures broad system-level tendencies or local deviations, and as such, the SHAP importance of a Gaussian process is aligned with the random forest results of Fig. 14.
@@ 6 Conclusions and perspectives
The codes and data used to produce the results are open-source and available online: https://github.com/ANR-MIMICO/SMPT_XAI.
@@ Appendix A Complementary Local Analysis of Variable Importance
In conclusion, SHAP values help explain model predictions, but their interpretation differs depending on the algorithm. Orange models, such as GP, reflect how each feature affects the predicted mean and uncertainty of the output, and SHAP highlights feature importance in shaping the underlying function rather than direct instance similarity. On the contrary, green models such as kNN make predictions based on the similarity between data points, using distances in feature space. SHAP values here indicate how much each feature contributes to determining the nearest neighbors and shaping the final prediction. There is no clear response on what is a “really important” variable, except that it depends on whether the interest of the practitioner is more in the global trends or in the potential local effects. Moreover, studying a complex system through a surrogate model is a strongly impacting method to gather information, as most of the resulting information comes from the structure of the surrogate model itself instead of coming from the complex system.